OpenStudy (anonymous):

please help n show work -7*(-10/27)*9/16

OpenStudy (anonymous):

here -7(-10/27)*9/16 first of all multiply -7 with -10/27 u would get : 70/27 * 9/16 (70*9)/(27*16) cut 9 with 27 since 9*3 = 27 so 3 would come in denominator on the pllace of 27 70/(3*16) 70/48 = 2(35)/2(48) = 35/24 = 1 whole 11/24

OpenStudy (anonymous):

did u understand this

OpenStudy (anonymous):

yes thank u so much

OpenStudy (anonymous):

remember here 1 whole 11/24 is written in mixed fraction

OpenStudy (anonymous):

no problem